The resistance of 120 meters of wire is 16Ω. What is the conductance of the wire from one end
to the middle?
1
Expert's answer
2021-09-17T15:32:56-0400

Conductance of a wire is inversely related to its resistance.

Resistance = 16 ohms

Conductance "= \frac{1}{16}\; ohms = 0.0625 \; mhos"

Per unit length conductance "= \frac{1}{16 \times 120} = 0.0005208 \; mho/m"

The conductance of the wire from one end to the middle "= 0.0005208 \times \frac{120}{2}"

"= 0.0005208 \times 60 = 0.03125 \;mhos"